# TAPO: Task-Referenced Adaptation for Prompt Optimization

### v2

@@ I Introduction
However, these methods face two primary limitations. First, current prompt evaluation techniques predominantly rely on a single metric, which hinders a comprehensive assessment. For example, PromptBreeder[6] and APE[2] employ a single similarity metric for fitness measurement, limiting their ability to logically improve tasks related to planning. Second, the lack of diverse metrics reduces their versatility, limiting their adaptability to multi-task optimization. For example, DATA [7] optimizes prompts for role-playing tasks, but exhibits limited scalability to a broader range of tasks. Similarly, while certain prompt strategies focused on machine translation [8] can improve language proficiency, their effectiveness in other essential language tasks, such as communication and reasoning, is diminished.
@@ III-A Experiment Settings
Baselines. We compare TAPO against the following baseline methods: (a) Zero-Shot CoT [15], which generates reasoning steps in a zero-shot manner; (b) APE [2], a method that initializes multiple prompt candidates from a base prompt and selects the best one based on development set performance; (c) PE2 [16], a two-step prompting approach that iteratively generates and refines candidate prompts through evaluation; and (d) PromptBreeder [6], a self-referential optimization framework that refines prompts by leveraging redescriptions to improve downstream task performance.
@@ III-C Task-Specific Prompt Performance (RQ2)
TAPO’s ability to tailor prompts for specific tasks significantly enhances performance across different domains, as illustrated in Table II. For math reasoning tasks like AddSub, TAPO emphasizes reasoning and computational steps, offering a customized approach that outperforms general methods like zero-shot CoT, which uses the generic prompt "Let’s think step by step". For translation tasks, TAPO excels by using a systematic framework for error categorization and iterative feedback, in contrast to the less structured methods of APE or PB. This task-specific design ensures that TAPO consistently delivers superior results by adapting the prompts to the unique requirements of each task, improving both clarity and performance in mathematical and translation error detection tasks.

### v3

@@ I Introduction
However, these methods face two primary limitations. First, current prompt evaluation techniques predominantly rely on a single metric, which hinders a comprehensive assessment. For example, PromptBreeder[6] and APE[2] employ a single similarity metric for fitness measurement, limiting their ability to logically improve tasks related to planning. Second, the lack of diverse metrics reduces their versatility, limiting their adaptability to multi-task optimization. For example, GATE [7] optimizes prompts for role-playing tasks, but exhibits limited scalability to a broader range of tasks. Similarly, while certain prompt strategies focused on machine translation [8] can improve language proficiency, their effectiveness in other essential language tasks, such as communication and reasoning, is diminished.
@@ III-A Experiment Settings
Baselines. We compare TAPO against the following baseline methods: (a) Zero-Shot CoT [15], which generates reasoning steps in a zero-shot manner; (b) APE [2], a method that initializes multiple prompt candidates from a base prompt and selects the best one based on development set performance; (c) PE2 [16], a two-step prompting approach that iteratively generates and refines candidate prompts through evaluation; and (d) PromptBreeder (PB)[6], a self-referential optimization framework that refines prompts by leveraging redescriptions to improve downstream task performance.
@@ III-C Task-Specific Prompt Performance (RQ2)
TAPO’s ability to tailor prompts for specific tasks significantly enhances performance across different domains, as illustrated in Table II. For math reasoning tasks like AddSub, TAPO emphasizes reasoning and computational steps, offering a customized approach that outperforms general methods like zero-shot CoT, which uses the generic prompt "Let’s think step by step". For translation tasks, TAPO excels by using a systematic framework for error categorization and iterative feedback, in contrast to the less structured methods of APE or PromptBreeder. This task-specific design ensures that TAPO consistently delivers superior results by adapting the prompts to the unique requirements of each task, improving both clarity and performance in mathematical and translation error detection tasks.